A concrete pipe, 2.44 meters long has for its outside diameter 533 mm. and inside diameter of 425 mm. Find the amount of concrete used in manufacturing 1,000 concrete pipes.
Answer by Cromlix(4381)

Hi there,
2.44 meters = 244 centimeters
533 millimeters = 53.3 centimeters
425 millimeters = 42.5 centimeters
Using outside radius (53.3/2 = 26.65 centimeters)
Volume = Pi x radius^2 x height(length)
Volume = Pi x 26.25^2 x 244 = 528199.8998 cm^3
Using inside radius (42.5/2 = 21.25)
Volume = Pi x radius^2 x height(length)
Volume = Pi x 21.25^2 x 244 = 346144.6056 cm^3
528199.8998 - 346144.6056
= 182055.2942 cm^3 This is the thickness of the pipe
The amount of concrete to make 1 pipe.
= 182.055 litres
1000 pipes = 182.055 x 1000
182055 litres of concrete.
Hope this helps :-)